OpenStudy (anonymous):

PRE-CALC QUESTION, HELP PLEASE 1.) Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

and this one too

rishavraj (rishavraj):

see if z = a + ib so \[z = \left| z \right| (\cos \theta + i \sin \theta)\]

rishavraj (rishavraj):

in the question a = -1 and \[b = 3\sqrt{3}\] so \[\left| z \right| = \sqrt{(-3)^2 + (3\sqrt{3})^2} \]

rishavraj (rishavraj):

*a = -3

OpenStudy (anonymous):

but what answer is it @rishavraj

rishavraj (rishavraj):

B

rishavraj (rishavraj):

@em2000 got it??

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

@rishavraj what about question 1